OpenStudy (anonymous):

Currently, Sam's coach is three times older than Sam. When Sam started soccer 5 years ago, his coach was five times his age. How old is Sam now? Answer 15 7 10 5

OpenStudy (mr.math):

You want to try?

OpenStudy (anonymous):

Ive been trying.

OpenStudy (anonymous):

what is it ten?

OpenStudy (mr.math):

What did you get?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

or is it 15

OpenStudy (mr.math):

Lets say that the Sam's age is x, and his coach is y, we can say that y=3x. But 5 years ago the coach was 5 times Sam's age, which means y-5=5(x-5). Now solve this equation by first substituting y=3x to get 3x-5=5x-25, which implies 2x=20 or x=10.

OpenStudy (mr.math):

So you're correct!